# AgentIR: Reasoning-Aware Retrival for Deep Research Agents

### 04 Mar 2026 v1

@@ 1 Introduction
We propose Reasoning-Aware Retrieval, a new retrieval paradigm that exploits this observation: instead of embedding only the agent’s issued query, Reasoning-Aware Retrieval jointly embeds the reasoning trace, learning a retriever that leverages the rich intent and contextual information expressed in agent reasoning. Further, to address the lack of retriever training data for agent-issued sub-queries in Deep Research, we introduce DR-Synth, a data synthesis method that transforms standard QA datasets such as WebShaper (Tao et al., 2025) into (agent sub-query, relevance) pairs tailored for Deep Research agent retrieval.
@@ 2 Related Work
A notable feature of Deep Research agents is their ability to interleave explicit reasoning and retrieval in solving complex tasks, which Reasoning-Aware Retrieval learns to leverage. In parallel, retrievers such as ReasonIR (Shao et al., 2025) have emerged to tackle reasoning-intensive tasks without an agent. We emphasis that Reasoning-Aware Retrieval addresses a fundamentally different objective: rather than expecting the retriever to solve a complex task in a single-turn, we focus on collaborative retrieval with an agent to resolve complex tasks over multiple turns.
@@ 4.2 Evaluation Setup and Metrics
To assess retrievers in end-to-end Deep Research, we pair them with three open-weight Deep Research agents: Tongyi-DeepResearch (Tongyi-DR) (Tongyi DeepResearch et al., 2025), gpt-oss-120B [^1] https://openai.com/index/introducing-gpt-oss/, and GLM-4.7222https://z.ai/blog/glm-4.7, evaluating them on BrowseComp-Plus (Chen et al., 2025), a benchmark featuring complex multi-hop queries requiring 20+ searches. Following its official evaluation, we give each agent a “search” tool that retrieves top-5 document snippets truncated to 512 tokens. Tongyi-DR is also trained to use a “visit” tool that opens a full document in addition to the “search” tool. For fair comparison, we record its accuracy both with and without this tool.
@@ 4.3 Baselines
As discussed in Section 3.2, Reasoning-Aware Retrieval’s method of embedding agent reasoning traces relates to past query expansion methods like HyDE. To this end, we compare with Reason-Rewriter [^4] https://huggingface.co/cfli/reasoner-rewriter-qwen2.5-7b-0821 + Reason-Embed-8B (Liu et al., 2025) [^5] https://huggingface.co/hanhainebula/reason-embed-qwen3-8b-0928, a finetuned HyDE-style query expander, paired with its dedicated retriever, shown to perform well on reasoning-intensive tasks. Further, we compare with Agentic-R (Liu et al., 2026), a concurrent work that also trains retrievers specialized for Deep Research agents, where they expand the agent’s query $q_{t}$ by prepending the global query $Q$, following the notation in Section 3.3.
@@
To provide a strong reference, we also evaluate listwise reranking (Sun et al., 2023), a computationally reranking method, applied to the first-stage Qwen3-Embedding-4B retriever. Specifically, we use Qwen3-8B (Zhang et al., 2025) to rerank the top-20 retrieved documents.
@@ Appendix B AgentIR-4B Training Details
  • Positives. For each positive document $p\in P$, we scrape its URL using Selenium  [^8] https://www.selenium.dev/documentation, and parse the content using Trafilatura (Barbaresi, 2021).

### 05 Mar 2026 v2

# AgentIR: Reasoning-Aware Retrieval for Deep Research Agents

@@ 1 Introduction
We propose Reasoning-Aware Retrieval, a new retrieval paradigm that exploits this observation: instead of embedding only the agent’s issued query, we jointly embed the reasoning trace, learning a retriever that leverages the rich intent and contextual information expressed in agent reasoning. Further, to address the lack of retriever training data for agent-issued sub-queries in Deep Research, we introduce DR-Synth, a data synthesis method that transforms standard QA datasets such as WebShaper (Tao et al., 2025) into (agent sub-query, relevance) pairs tailored for Deep Research agent retrieval.
@@ 2 Related Work
A notable feature of Deep Research agents is their ability to interleave explicit reasoning and retrieval in solving complex tasks, which we learn to leverage. In parallel, retrievers such as ReasonIR (Shao et al., 2025) and RaDeR (Das et al., 2025) have emerged to tackle reasoning-intensive tasks without an agent. We emphasize that Reasoning-Aware Retrieval addresses a fundamentally different objective: rather than expecting the retriever to solve a complex task in a single-turn, we focus on collaborative retrieval with an agent to resolve complex tasks over multiple turns.
@@ 4.2 Evaluation Setup and Metrics
To assess retrievers in end-to-end Deep Research, we pair them with three open-weight Deep Research agents: Tongyi-DeepResearch (Tongyi-DR) (Tongyi DeepResearch et al., 2025), gpt-oss-120B, [^1] https://openai.com/index/introducing-gpt-oss/ and GLM-4.7, [^2] https://z.ai/blog/glm-4.7 evaluating them on BrowseComp-Plus (Chen et al., 2025), a benchmark featuring complex multi-hop queries requiring 20+ searches. Following its official evaluation, we give each agent a “search” tool that retrieves top-5 document snippets truncated to 512 tokens. Tongyi-DR is also trained to use a “visit” tool that opens a full document in addition to the “search” tool. For fair comparison, we record its accuracy both with and without this tool.
@@ 4.3 Baselines
As discussed in Section 3.2, Reasoning-Aware Retrieval’s method of embedding agent reasoning traces relates to past query expansion methods like HyDE. To this end, we compare with Reason-Rewriter [^4] https://huggingface.co/cfli/reasoner-rewriter-qwen2.5-7b-0821 + Reason-Embed-8B (Liu et al., 2025), [^5] https://huggingface.co/hanhainebula/reason-embed-qwen3-8b-0928 a finetuned HyDE-style query expander, paired with its dedicated retriever, shown to perform well on reasoning-intensive tasks. Further, we compare with Agentic-R (Liu et al., 2026), a concurrent work that also trains retrievers specialized for Deep Research agents, where they expand the agent’s query $q_{t}$ by prepending the global query $Q$, following the notation in Section 3.3.
@@
To provide a strong reference, we also evaluate listwise reranking (Sun et al., 2023), a computationally reranking method, applied to the first-stage Qwen3-Embedding-4B retriever. Specifically, we use Qwen3-8B (Yang et al., 2025) to rerank the top-20 retrieved documents.
@@ Appendix B AgentIR-4B Training Details
  • Positives. For each positive document $p\in P$, we scrape its URL using Selenium, [^8] https://www.selenium.dev/documentation and parse the content using Trafilatura (Barbaresi, 2021).